OpenStudy (anonymous):

2(x - 4) + 6(3 - x) = 7.

OpenStudy (anonymous):

\[2x-8+18-6x=7\]

OpenStudy (anonymous):

\[-4x+10=7\]

OpenStudy (anonymous):

wont x = -7

OpenStudy (anonymous):

\[-4x=-3\]

OpenStudy (anonymous):

x = 3/4